OpenStudy (help!!!!):

Mitch drives for 30min at 50mph. He then drives for an additional 45 min at 60 mph. How many miles did he drive altogether? a 60mile b80mile cc50miles d65miles e70miles. &how did u find it out, is there a formula for it?

OpenStudy (anonymous):

is doesnt make sense

OpenStudy (kainui):

So lets think about it. If you 50 miles per hour in 1 hour, how far have you travelled? It makes sense that you would have traveled 50 miles right? So the formula looks like: distance=rate*time or simpler: d=rt Now if you're travelling for only 30 minutes, well since your rate is in hours you have to convert over the time to hours or the rate to minutes. I would suggest converting the time to hours since it seems simpler. You already know that 30 minutes is half of an hour, right? So to convert from minutes to hours you have to divide by 60. \[\frac{ 30 minutes }{ 60 minutes/ hour }=\frac{ 1 }{ 2 } hour\] see if you can figure it out. It might take a few minutes to think about all this and "get it" but once you do you'll be able to do all these types of problems by just reasoning them out like this which will really help on tests!

OpenStudy (anonymous):

It's a lie, Mitch doesn't have a licence!

OpenStudy (anonymous):

what

OpenStudy (anonymous):

30 mins @ 50 miles an hour is 25 miles. i.e. 50 miles in half an hour 50 * 1/2 = 50/2 = 25 miles. Now you cannae work out the resst.

OpenStudy (anonymous):

@ 60 mph Mitch is doing a mile a minute does that make sense?

OpenStudy (anonymous):

I'd gues at about 60+ milesz

OpenStudy (anonymous):

r t = d \[\left(50*\frac{30}{60}\right)+\left(60*\frac{45}{60}\right)=70 \text{ miles}\]